# SUG-Occ: An Explicit Semantics and Uncertainty Guided Sparse Learning Framework for Real-Time 3D Occupancy Prediction

### v1

@@ I Introduction
The 3D semantic occupancy prediction has emerged as an fundamental perception task for scene understanding in autonomous driving, thanks to its ability to jointly estimate voxel-level geometric and semantic states in 3D space [45]. Differing from conventional perception paradigms like image segmentation [16, 8], bounding-box detection [47, 22], and Bird-Eye-View (BEV) map [28, 33], 3D semantic occupancy prediction produce dense and spatially consistent 3D scene representations which is beneficial for downstream tasks such as motion planning and other safety-critical decision-making in real-time autonomous systems.
@@ III-F Training Objectives
TABLE I: Overall 3D semantic occupancy prediction result on SemanticKITTI validation set. For each method, we report its venue, image backbone, parameter counts, geometry IoU, mean semantic IoU and FPS. * represents these methods are adapted for the image inputs, which are implemented and reported in MonoScene [5].
@@ IV-C Main Results
We compare the porposed method with several representative and state-of-the-art 3D semantic occupancy prediction methods on the SemanticKITTI dataset, including MonoScene [5], TPVFormer [20], VoxFormer [26], OccFormer [49], SparseOcc [41] and ProtoOcc [21]. For performance comparison, the reported results of these methods are taken directly derived from their original papers. To ensure a fair evaluation of efficiency, we re-evaluate all models under their official configurations in a unified environment. Models with publicly available checkpoints are evaluated through direct inference. For methods without released weights, we retrain them from scratch following their reported configurations.
Table I presents an overall comparison on the SemanticKITTI validation set, including the image backbone, number of parameters, geometric IoU, mean IoU (mIoU) over all semantic classes and inference speed measured in FPS. As shown in the table, although using a relatively lightweight ResNet-50 based image backbone, our method achieve the highest mIoU and FPS among all compared approaches. In particular, compared with ProtoOcc, the proposed method improves mIoU by $7.42\%$ while achieving a $57.81\%$ increase in FPS. Note that the parameter counts and FPS of MonoScene, TPVFormer and VoxFormer are not reported as their official implementations rely on legacy versions of PyTorch and CUDA that are incompatible with our evaluation environment.

### v2

@@ I Introduction
The 3D semantic occupancy prediction has emerged as an fundamental perception task for scene understanding in autonomous driving, thanks to its ability to jointly estimate voxel-level geometric and semantic states in 3D space [47]. Differing from conventional perception paradigms like image segmentation [16, 8], bounding-box detection [49, 22], and Bird-Eye-View (BEV) map [28, 33], 3D semantic occupancy prediction produce dense and spatially consistent 3D scene representations which is beneficial for downstream tasks such as motion planning [46] and other safety-critical decision-making [36] in real-time autonomous systems.
@@ III-F Training Objectives
TABLE I: Overall 3D semantic occupancy prediction result on SemanticKITTI validation set. For each method, we report its venue, image backbone, parameter counts, geometry IoU, mean semantic IoU and FPS. * represents these methods are adapted for the image inputs, which are implemented and reported in MonoScene [5]. - indicates that the corresponding information is not applicable or the official implementations are incompatible with our environment.
@@ IV-C Main Results
We compare the proposed method with several representative and state-of-the-art 3D semantic occupancy prediction methods on the SemanticKITTI dataset, including MonoScene [5], TPVFormer [20], VoxFormer [26], OccFormer [51], SparseOcc [42] and ProtoOcc [21], as well as several earlier methods such as LMSCNet [38], 3DSketch [5], AICNet [9], and JS3C-Net[48], whose results are reproduced and reported by MonoScene. For performance comparison, the reported results of these methods are taken directly derived from their original papers. To ensure a fair evaluation of efficiency, we re-evaluate all models under their official configurations in a unified environment. Models with publicly available checkpoints are evaluated through direct inference. For methods without released weights, we retrain them from scratch following their reported configurations.
Table I presents an overall comparison on the SemanticKITTI validation set, including the image backbone, number of parameters, geometric IoU, mean IoU (mIoU) over all semantic classes and inference speed measured in FPS. As shown in the table, although using a relatively lightweight ResNet-50 based image backbone, our method achieves the highest mIoU and FPS among all compared approaches. In particular, compared with the state-of-the-art method ProtoOcc, the proposed method improves mIoU by $7.42\%$ while achieving a $57.81\%$ increase in FPS. Note that the parameter counts and FPS of MonoScene, TPVFormer and VoxFormer are not reported as their official implementations rely on legacy versions of PyTorch and CUDA that are incompatible with our evaluation environment, while the image backbones of those earlier methods are not applicable, which are both marked as - in the table.